# Revisiting Zeroth-Order Optimization for Memory-Efficient LLM Fine-Tuning: A Benchmark

### v2

@@ 1 Introduction
To address (Q), our work introduces several key innovations compared to the most relevant work (Malladi et al., 2023). We explore a broader range of ZO optimization methods beyond ZO-SGD and examine various task and model types, and evaluation metrics. We conduct a detailed comparative analysis of different ZO optimization methods, shedding light on the often-overlooked forward gradient method (Ren et al., 2022) and other ZO optimization techniques in LLM fine-tuning. This benchmarking study helps reveal the pros and cons of these methods in accuracy and efficiency. Extended from the gained insights, we propose to further improve ZO optimization-based LLM fine-tuning using techniques of block-wise descent, hybrid ZO and FO training, and gradient sparsity.
In summary, our key contributions are listed below.
@@ 2 Related Work
Zeroth-order optimization. Zeroth-order (ZO) optimization is a technique that uses finite differences to estimate gradients. Such algorithms utilize function value oracle only, yet share similar algorithm structure with first-order (FO) gradient-based counterpart methods. ZO optimization usually enjoys provable (dimension-dependent) convergence guarantees, as discussed in various works (Flaxman et al., 2005; Nesterov & Spokoiny, 2017; Duchi et al., 2015; Ghadimi & Lan, 2013). These methods have draw considerable attention due to its effectiveness in a wide range of modern machine learning (ML) challenges (Liu et al., 2020), including the adversarial attack and defense (Chen et al., 2017; Tu et al., 2019; Ye et al., 2018; Ilyas et al., 2018; Zhang et al., 2022b; Verma et al., 2023; Zhao et al., 2019; Hogan & Kailkhura, 2018; Shu et al., 2022), model-agnostic contrastive explanations (Dhurandhar et al., 2019), enhancing transfer learning through visual prompting (Tsai et al., 2020), computational graph unrolling (Vicol et al., 2023), and optimizing automated ML processes (Gu et al., 2021a; Wang et al., 2022). Beyond standard ML, it finds application in policy search in reinforcement learning (Vemula et al., 2019), network resource management (Liu et al., 2018), ML-based optimization of scientific workflows (Hoffman et al., 2022; Tsaknakis et al., 2022; Chen et al., 2024), and on-chip learning enhancements (Gu et al., 2021b).
@@
BP-free training for large models. Training large models, especially LLMs, is memory-consuming due to the involved large computation graphs for BP (Ren et al., 2021; Kim et al., 2023). Thus BP-free methods have become a recent focus in the deep learning (DL) community. Forward gradient learning (Baydin et al., 2022; Ren et al., 2022; Silver et al., 2021; Belouze, 2022), built upon the forward-mode automatic differentiation (AD), provides an alternative to ZO optimization for BP-free training. Different from ZO optimization, it relies on the forward-mode AD to calculate a forward (directional) gradient. However, one main limitation of the forward gradient is its requirement of full access to the AD software and the deep model, making it less memory-efficient than ZO optimization and impractical for tackling black-box problems (Chen et al., 2024). The specifications of BP-free methods include greedy layer-wise learning (Nøkland & Eidnes, 2019), input-weight alignment (Boopathy & Fiete, 2022), forward-forward algorithm (Hinton, 2022), synthetic gradients (Jaderberg et al., 2017), BBT/BBTv2 evolutionary algorithms (Sun et al., 2022a, b), gradient guessing using special low dimensional structure of neural networks (Singhal et al., 2023) and other black-box methods which optimize the prompts (Prasad et al., 2023; Deng et al., 2022; Chai et al., 2022). Many of these algorithms are also motivated through the lens of seeking DL’s biological interpretation.
Applying ZO optimization to fine-tune pre-trained LLMs is particularly intriguing because it combines the advantages of being BP-free and utilizing pre-training. This enables the scalability of ZO optimization to large-scale LLMs while maintaining memory efficiency. MeZO (Malladi et al., 2023) introduced a scalable ZO-SGD algorithm to efficiently fine-tune LLMs with up to 60 billion parameters, achieving competitive results compared to first-order optimization methods and structured fine-tuning approaches like LoRA. They also provided theoretical insights into why ZO methods can be effective for LLMs. This opens the gateway for efficient BP-free LLM fine-tuning and largely motivates our ZO benchmark study.
@@ 3 Reviewing ZO Optimization and Beyond
As a byproduct of connecting RGE to (1), we obtain the directional derivative-based gradient estimate, $\nabla f(\mathbf{x})\approx f^{\prime}(\mathbf{x},\mathbf{u})\mathbf{u}$, which is known as the forward gradient (Forward-Grad) (Baydin et al., 2022; Ren et al., 2022). Different from RGE that relies solely on the finite difference of function values, Forward-Grad requires the use of forward mode automatic differentiation (AD) but eliminates the need for the backward evaluation in the implementation of deep model fine-tuning or training. In other words, Forward-Grad is BP-free and can serve as another alternative gradient estimation method that improves the memory efficiency of LLM fine-tuning. We stress that Forward-Grad is a possibly overlooked BP-free optimizer. Given its unbiasedness as shown in (1), it could serve as an upper performance bound for ZO optimization in theory.
@@
where $\mathbf{x}_{t}$ denotes the updated solution at the $t$th iteration, $\eta_{t}>0$ is the learning rate, and $h(\cdot)$ is a certain descent direction post-processing operation. In (2), we omit the inclusion of the stochastic mini-batch for empirical risk minimization for ease of presentation. For instance, ZO-SGD can be expressed as (2) when $h(\hat{\nabla}f(\mathbf{x}))=\hat{\nabla}f(\mathbf{x})$. Similarly, ZO-SGD-Sign can be derived if $h(\hat{\nabla}f(\mathbf{x}))=\mathrm{sign}(\hat{\nabla}f(\mathbf{x}))$, where $\mathrm{sign}(\cdot)$ represents element-wise sign operation. Another example is ZO-SGD-Cons by setting $h(\hat{\nabla}f(\mathbf{x}))=\operatorname*{arg\,min}_{\mathbf{g}\in\{\mathbf{0},-\hat{\nabla}f(\mathbf{x}),\hat{\nabla}f(\mathbf{x})\}}f(\mathbf{x}_{t}+\eta_{t}\mathbf{g})$. We refer readers to Appx. B for more algorithmic details of (2) as applied to the ZO optimization approaches we covered.
@@
As a warm-up experiment, Tab. 1 empirically justifies the importance of task alignment when applying ZO optimization to LLM fine-tuning on the simple binary classification task by comparing scenarios with and without the use of pre-defined prompts to achieve task alignment. We fine-tune the entire Roberta-Large (Liu et al., 2019b) model on SST2 (Socher et al., 2013) and RTE (Wang et al., 2019) datasets with two selected ZO methods: ZO-SGD (i.e., MeZO in (Malladi et al., 2023)) and ZO-Adam. We compare their performance with that of the FO method (FO-SGD). The task alignment is achieved with the template <CLS>SENTENCE. It was [terrible|great].<SEP> for SST dataset and another template <CLS>SENTENCE1? [Yes|No], SENTENCE2.<SEP> for RTE. As we can see, without prompt-based text alignment, there is a substantial performance drop across ZO fine-tuning methods. Both ZO-SGD and ZO-Adam yield about $10\%$ and $8\%$ accuracy degradation on SST2 and RTE, respectively. In contrast, FO-SGD suffers less from the absence of task alignment. This suggests that the task alignment is particularly beneficial to ZO LLM fine-tuning. It is also worth noting that crafting effective prompts for task alignment can be non-trivial, as prompt design is context-dependent and can affect the fine-tuning performance. In this work, we follow (Gao et al., 2020) and (Malladi et al., 2023) to align the fine-tuning tasks to the pretrained ones.
4 Benchmarking ZO Optimization for LLM Fine-Tuning
In this section, we delve into the empirical performance of ZO optimization in LLM fine-tuning. Our benchmarking effort includes evaluating accuracy and efficiency, accounting for different downstream task complexities (ranging from simple classification to more complicated reasoning task), and considering various language model types.
@@ 4.1 Benchmark Setups
LLM fine-tuning tasks, schemes, and models. We begin by introducing the tasks and the fine-tuning schemes. We focus on three tasks, considering their complexity from low to high, which include (1) the simplest binary classification task, Stanford Sentiment Treebank v2 (SST2) (Socher et al., 2013), (2) the question-answering task, Choice Of Plausible Alternatives (COPA) (Roemmele et al., 2011), and (3) the commonsense reasoning task, WinoGrande (Sakaguchi et al., 2021). When evaluating memory efficiency, we also consider the task of multi-sentence reading comprehension (MultiRC) (Khashabi et al., 2018). For LLM fine-tuning on these tasks, we explore four parameter-efficient fine-tuning (PEFT) schemes: full-tuning (FT) that fine-tunes the entire pre-trained model, low-rank adaptation (LoRA) by imposing low-rank weight perturbations (Hu et al., 2021a), prefix-tuning (Prefix) by appending learnable parameters to token embedding (Li & Liang, 2021), and prompt-tuning (Prompt) (Liu et al., 2022) by introducing a series of learnable tokens attached to the input to adapt the fixed model to downstream tasks. We refer readers to Appx. A for details. Furthermore, we incorporate several representative language models, including Roberta-Large (Liu et al., 2019b), OPT (Zhang et al., 2022a), LLaMA2 (Touvron et al., 2023), Vicuna (Zheng et al., 2023), and Mistral (Jiang et al., 2023).
@@
Setup and implementation of ZO optimization. To train the previously mentioned LLM fine-tuners, we utilize the ZO optimization methods introduced in Sec. 3. These include ZO-SGD (i.e. MeZO (Malladi et al., 2023)), ZO-SGD-Sign, ZO-SGD-MMT, ZO-SGD-Cons, and ZO-Adam. For comparison, we also present the performance of Forward-Grad, which relies on the forward mode auto-differentiation rather than BP. We also provide the performance of two FO optimizers: (FO-)SGD and (FO-)Adam. Before applying the aforementioned optimizers to the LLM fine-tuning tasks, we follow (Gao et al., 2020; Malladi et al., 2023) to align the fine-tuning task format with the token or sentence prediction-based pre-training task, as demonstrated in Tab. 1. We run ZO (or BP-free) optimizers and FO optimizers for 20000 and 625 iterations respectively, as outlined in (2). Note that ZO optimization takes a longer convergence time as shown in (Liu et al., 2020). When implementing (RGE), unless specified otherwise, we set the query budget per gradient estimation to $q=1$. We determine the values of other hyperparameters, such as the smoothing parameter and learning rate, through a grid search for each method. Following (Malladi et al., 2023), if not specified otherwise, we adopt the mixed precision (MP) training for FO optimizers, where the model is loaded with the full precision (i.e., 32 bits) but training is carried out in half precision with 16 bits (FP16). For ZO optimizers, half-precision training is adopted throughout the model loading and training. We also remark that neither MP nor FP16 can be applied to Forward-Grad. We refer readers for more details to Appx. C.
@@ 4.2 Experiment Results
Third, The performance of ZO methods exhibits high variance, as evidenced by the fluctuating relative rankings across different experimental scenarios, although extensive hyper-parameter search efforts have been made. For example, the effectiveness of ZO-Adam degrades dramatically in the (OPT-1.3B, Prompt) setting. In addition, the MeZO method (i.e., ZO-SGD) used in (Malladi et al., 2023) does not always emerge as the top-performing ZO optimizer for LLM fine-tuning across various settings. This variance is not surprising and could be largely attributed to the high variance of RGE (Nesterov & Spokoiny, 2017; Duchi et al., 2015).
@@
First, compared to the previous results, the performance gap among different ZO methods are much enlarged. In the meantime, the performance gap between FO and ZO methods are also widened. For example, in the experiments with WinoGrande, the FO methods (FO-SGD and FO-Adam) outperform all the other ZO methods by a large margin. This observation shows the scalability bottleneck intrinsic to ZO methods, when dealing with larger models and/or more complex tasks.
Second, certain ZO methods exhibit exceptional stability across varied conditions: Despite a general trend towards variability, specific ZO methods, e.g., ZO-Adam and ZO-SGD-MMT, demonstrate stability in their performance. The reason for this could be that these algorithms integrate variance-reduced optimization techniques (such as momentum and adaptive learning rate) to ZO optimization and become more adaptable and resilient to the variances of ZO gradient estimation (Chen et al., 2019).
Third, the LoRA tuning method is consistently robust when paired with various ZO algorithms. This resilience across different ZO methods suggests that LoRA’s mechanism is inherently more adaptable to the variations in ZO optimization strategies, providing a stable and reliable tuning approach in diverse settings. We will peer into the performance of LoRA below.
@@
In Tab. 3, we present how different ZO methods perform on (LoRA, COPA) and (LoRA, WinoGrande) across a range of widely-used LLM families. For ease of computation, we focus on a subset of ZO optimization methods, including ZO-SGD, ZO-SGD-CONS, and ZO-Adam. As we can see, in some scenarios with the COPA dataset, some BP-free methods exhibit effectiveness comparable to, or even superior to, that of FO methods (FO-SGD and FO-Adam). For example, Forward-Grad and ZO-Adam outperform the best FO method on model OPT-13B and Vicuna-7B. Conversely, for the more difficult task WinoGrande, a performance gap of $5\%\sim 6\%$ between FO and ZO methods still exists across different models.
Figure 2: Memory comparison between FO-SGD and ZO-SGD full fine-tuning across various sequence lengths with a fixed effective batch size of $2$. Memory evaluation was conducted using synthetic text generated from random sequences of the specified shapes. For shorter sequences (i.e., $<700$), the memory usage of FO-SGD remains relatively stable since the memory consumption for storing gradients during BP surpasses that needed for activations.
Efficiency analyses. In Tab. 4, we present a comparison of the efficiency performance of various ZO/FO optimizers when fine-tuning the full OPT-13B model on the MultiRC dataset with a batch size $4$. We evaluate the efficiency in the following dimensions: memory cost (in GB), the consumption of GPU resources (number of GPUs), and runtime cost per optimization iteration (in seconds). All the experiments are carried out in the same environment. Several key observations can be made below.
First, almost all ZO methods (except ZO-Adam) demonstrate comparable levels of efficiency, requiring only a single GPU (A100) for LLM fine-tuning. This observation is expected as ZO methods entail relatively straightforward optimization steps, primarily based on function evaluations, as depicted in RGE. Among the examined ZO optimization methods, ZO-Adam incurs higher memory costs due to its algorithmic complexity. Second, in comparison to FO methods, ZO optimization reduces runtime costs per iteration, e.g., by approximately $33.3\%$ for ZO-SGD compared to FO-SGD. Third, Forward-Grad appears to be the point at which ZO optimization methods lose their memory efficiency advantage over FO methods. Additionally, we note that the substantial runtime cost of Forward-Grad compared to FO optimizers (FO-SGD and FO-Adam) is possibly attributed to its incompatibility with MP and FP16.
Furthermore, we examine the memory cost of LLM fine-tuning vs. the input sequence length. In Fig. 2, we compare the memory efficiency between ZO-SGD and FO-SGD across various sequence lengths (i.e. the token number per sample). We employed synthetic texts generated from random sequences with specified shapes. In all experiments, an effective batch size of $2$ was maintained. The memory consumption of ZO-SGD maintains a consistent level, whereas FO-SGD begins to demand significantly more memory as the sequence length extends. This trend is amplified particularly for longer sequences (e.g., exceeding $700$ as depicted in Fig. 2), where the memory allocated for activations overwhelms that required for gradient storage. We provide a theoretical analysis of this phenomenon in Appx C.
Table 4: The peak memory cost (in GB), the required GPU resources, and the runtime cost (in seconds) of each optimizer when fine-tuning the full OPT-13B model on MultiRC with an averaged 400 context length. The order of included optimizers is ranked based on the memory cost. The per-iteration runtime in seconds (s) is averaged over 100 iterations. Notably, Forward-Grad is marked by ***, indicating its incompatibility with efficiency-enhancing techniques such as MP (mixed-precision training) and FP16 (half-precision training).
@@
ZO-SGD $\mathbf{29}$ GB $\mathbf{1}\times$A100 $\mathbf{1.8}$s
ZO-SGD-Cons $\mathbf{29}$ GB $\mathbf{1}\times$A100 $4.2$s
ZO-SGD-Sign $\mathbf{29}$ GB $\mathbf{1}\times$A100 $\mathbf{1.8}$s
ZO-SGD-MMT $53$ GB $\mathbf{1}\times$A100 $\mathbf{1.8}$s
ZO-Adam $80$ GB 2×2$\times$2 ×A100 $1.9$s
$*{}^{*}$ $138$ GB 2×2$\times$2 ×A100 $19.8$s
FO-SGD $161$ GB 3×3$\times$3 ×A100 $2.7$s
FO-Adam $257$ GB 4×4$\times$4 ×A100 $2.8$s
Ablation study on query budget. Recall from (1) that Forward-Grad provides us with an unbiased gradient estimate with respect to the FO gradient, in contrast to the function value-based biased ZO gradient estimate. However, in the above experiments, we have not observed a significant advantage brought by Forward-Grad. We hypothesize that this is due to the smallest query budget $q=1$ we used, which, despite its query efficiency, may not fully showcase the unbiased benefit.

### v3

@@ 1 Introduction
To address (Q), our work introduces several key innovations compared to the most relevant work (Malladi et al., 2023). We explore a broader range of ZO optimization methods beyond ZO-SGD and examine various task and model types, as well as evaluation metrics. We conduct a detailed comparative analysis of different ZO optimization methods, shedding light on the often-overlooked forward gradient method (Ren et al., 2022) and other ZO optimization techniques in LLM fine-tuning. This benchmarking study helps reveal the pros and cons of these methods in accuracy and efficiency. Extended from the gained insights, we propose to further improve ZO optimization-based LLM fine-tuning using techniques of block-wise descent, hybrid ZO and FO training, and gradient sparsity. In summary, our key contributions are listed below.
@@ 2 Related Work
Zeroth-order optimization. Zeroth-order (ZO) optimization is a technique that uses finite differences to estimate gradients. Such algorithms utilize function value oracle only, yet share a similar algorithm structure with first-order (FO) gradient-based counterpart methods. ZO optimization usually enjoys provable (dimension-dependent) convergence guarantees, as discussed in various works (Flaxman et al., 2005; Nesterov & Spokoiny, 2017; Duchi et al., 2015; Ghadimi & Lan, 2013). These methods have drawn considerable attention due to their effectiveness in a wide range of modern machine learning (ML) challenges (Liu et al., 2020), including the adversarial attack and defense (Chen et al., 2017; Tu et al., 2019; Ye et al., 2018; Ilyas et al., 2018; Zhang et al., 2022b; Verma et al., 2023; Zhao et al., 2019; Hogan & Kailkhura, 2018; Shu et al., 2022), model-agnostic contrastive explanations (Dhurandhar et al., 2019), enhancing transfer learning through visual prompting (Tsai et al., 2020), computational graph unrolling (Vicol et al., 2023), and optimizing automated ML processes (Gu et al., 2021a; Wang et al., 2022). Beyond standard ML, it finds application in policy search in reinforcement learning (Vemula et al., 2019), network resource management (Liu et al., 2018), ML-based optimization of scientific workflows (Hoffman et al., 2022; Tsaknakis et al., 2022; Chen et al., 2024), and on-chip learning enhancements (Gu et al., 2021b).
@@
BP-free training for large models. Training large models, especially LLMs, is memory-consuming due to the involved large computation graphs for BP (Ren et al., 2021; Kim et al., 2023). Thus, BP-free methods have recently become a focus in the deep learning (DL) community. Forward gradient learning (Baydin et al., 2022; Ren et al., 2022; Silver et al., 2021; Belouze, 2022), built upon the forward-mode automatic differentiation (AD), provides an alternative to ZO optimization for BP-free training. Unlike ZO optimization, it relies on the forward-mode AD to calculate a forward (directional) gradient. However, one main limitation of the forward gradient is its requirement of full access to the AD software and the deep model, making it less memory-efficient than ZO optimization and impractical for tackling black-box problems (Chen et al., 2024). The specifications of BP-free methods include greedy layer-wise learning (Nøkland & Eidnes, 2019), input-weight alignment (Boopathy & Fiete, 2022), forward-forward algorithm (Hinton, 2022), synthetic gradients (Jaderberg et al., 2017), BBT/BBTv2 evolutionary algorithms (Sun et al., 2022a, b), gradient guessing using special low dimensional structure of neural networks (Singhal et al., 2023) and other black-box methods which optimize the prompts (Prasad et al., 2023; Deng et al., 2022; Chai et al., 2022). Many of these algorithms are also motivated by seeking DL’s biological interpretation.
Applying ZO optimization to fine-tune pre-trained LLMs is particularly intriguing because it combines the advantages of being BP-free and utilizing pre-training. This improves the scalability of ZO optimization to large-scale LLMs while maintaining memory efficiency. MeZO (Malladi et al., 2023) introduced a ZO-SGD algorithm to fine-tune LLMs with up to 60 billion parameters, very competitive compared to first-order optimization methods and structured fine-tuning approaches like LoRA. They also provided theoretical insights into why ZO methods can be effective for LLMs. This opens the gateway for efficient BP-free LLM fine-tuning and largely motivates our ZO benchmark study.
@@ 3 Reviewing ZO Optimization and Beyond
As a byproduct of connecting RGE to (1), we obtain the directional derivative-based gradient estimate, $\nabla f(\mathbf{x})\approx f^{\prime}(\mathbf{x},\mathbf{u})\mathbf{u}$, which is known as the forward gradient (Forward-Grad) (Baydin et al., 2022; Ren et al., 2022). Different from RGE that relies solely on the finite difference of function values, Forward-Grad requires the use of forward mode automatic differentiation (AD) but eliminates the need for backward evaluation in the implementation of deep model fine-tuning or training. In other words, Forward-Grad is BP-free and can serve as another alternative gradient estimation method that improves the memory efficiency of LLM fine-tuning. We stress that Forward-Grad is a possibly overlooked BP-free optimizer. Given its unbiasedness as shown in (1), it could serve as an upper performance bound for ZO optimization in theory.
@@
where $\mathbf{x}_{t}$ denotes the updated solution at the $t$th iteration, $\eta_{t}>0$ is the learning rate, and $h(\cdot)$ is a certain descent direction post-processing operation. In (2), we omit the inclusion of the stochastic mini-batch for empirical risk minimization for ease of presentation. For instance, ZO-SGD can be expressed as (2) when $h(\hat{\nabla}f(\mathbf{x}))=\hat{\nabla}f(\mathbf{x})$. Similarly, ZO-SGD-Sign can be derived if $h(\hat{\nabla}f(\mathbf{x}))=\mathrm{sign}(\hat{\nabla}f(\mathbf{x}))$, where $\mathrm{sign}(\cdot)$ represents element-wise sign operation. Another example is ZO-SGD-Cons by setting $h(\hat{\nabla}f(\mathbf{x}))=\operatorname*{arg\,min}_{\mathbf{g}\in\{\mathbf{0},-\hat{\nabla}f(\mathbf{x}),\hat{\nabla}f(\mathbf{x})\}}f(\mathbf{x}_{t}+\eta_{t}\mathbf{g})$. We refer readers to Appendix A for more algorithmic details of (2) as applied to the ZO optimization approaches.
@@
As a warm-up experiment, Tab. 1 empirically justifies the importance of task alignment when applying ZO optimization to LLM fine-tuning on the simple binary classification task by comparing scenarios with and without the use of pre-defined prompts to achieve task alignment. We fine-tune the entire Roberta-Large (Liu et al., 2019b) model on SST2 (Socher et al., 2013) and RTE (Wang et al., 2019) datasets with two selected ZO methods: ZO-SGD (i.e., MeZO in (Malladi et al., 2023)) and ZO-Adam. We compare their performance with that of the FO method (FO-SGD). The task alignment is achieved with the template <CLS>SENTENCE. It was [terrible|great].<SEP> for the SST dataset and another template <CLS>SENTENCE1? [Yes|No], SENTENCE2.<SEP> for RTE. As we can see, without prompt-based text alignment, there is a substantial performance drop across ZO fine-tuning methods. Both ZO-SGD and ZO-Adam yield about $10\%$ and $8\%$ accuracy degradation on SST2 and RTE, respectively. In contrast, FO-SGD suffers less from the absence of task alignment. This suggests that the task alignment is particularly beneficial to ZO LLM fine-tuning. It is also worth noting that crafting effective prompts for task alignment can be non-trivial, as prompt design is context-dependent and can affect the fine-tuning performance. In this work, we follow (Gao et al., 2020) and (Malladi et al., 2023) to align the fine-tuning tasks to the pretrained ones.
4 LLM Fine-Tuning Benchmarking
In this section, we delve into the empirical performance of ZO optimization in LLM fine-tuning. Our benchmarking effort includes evaluating accuracy and efficiency, accounting for different downstream task complexities (ranging from simple classification to more complicated reasoning tasks), and considering various language model types.
@@ 4.1 Benchmark Setups
LLM fine-tuning tasks, schemes, and models. We begin by introducing the tasks and the fine-tuning schemes. We focus on three tasks, considering their complexity from low to high, which include (1) the simplest binary classification task, Stanford Sentiment Treebank v2 (SST2) (Socher et al., 2013), (2) the question-answering task, Choice Of Plausible Alternatives (COPA) (Roemmele et al., 2011), (3) the commonsense reasoning task, WinoGrande (Sakaguchi et al., 2021), and (4) the multi-sentence reading comprehension (MultiRC) (Khashabi et al., 2018) (for efficiency evaluation only). For LLM fine-tuning on these tasks, we explore four parameter-efficient fine-tuning (PEFT) schemes: full-tuning (FT) that fine-tunes the entire pre-trained model, low-rank adaptation (LoRA) by imposing low-rank weight perturbations (Hu et al., 2021a), prefix-tuning (Prefix) by appending learnable parameters to token embedding (Li & Liang, 2021), and prompt-tuning (Prompt) (Liu et al., 2022) by introducing a series of learnable tokens attached to the input to adapt the fixed model to downstream tasks. We refer readers to Appx. B for details. Furthermore, we incorporate several representative language models, including Roberta-Large (Liu et al., 2019b), OPT (Zhang et al., 2022a), LLaMA2 (Touvron et al., 2023), Vicuna (Zheng et al., 2023), and Mistral (Jiang et al., 2023).
@@
Setup and implementation details. To train the previously mentioned LLM fine-tuners, we utilize the ZO optimization methods introduced in Sec. 3. These include ZO-SGD (i.e. MeZO (Malladi et al., 2023)), ZO-SGD-Sign, ZO-SGD-MMT, ZO-SGD-Cons, and ZO-Adam. For comparison, we also present the performance of Forward-Grad, which relies on the forward mode auto-differentiation rather than BP. We also provide the performance of two FO optimizers: (FO-)SGD and (FO-)Adam. Before applying the aforementioned optimizers to the LLM fine-tuning tasks, we follow (Gao et al., 2020; Malladi et al., 2023) to align the fine-tuning task format with the token or sentence prediction-based pre-training task, as demonstrated in Tab. 1. We run ZO (or BP-free) optimizers and FO optimizers for 20000 and 625 iterations respectively, as outlined in (2). Note that ZO optimization takes a longer convergence time as shown in (Liu et al., 2020). When implementing (RGE), unless specified otherwise, we set the query budget per gradient estimation to $q=1$. We determine the values of other hyperparameters, such as the smoothing parameter and learning rate, through a grid search for each method. Unless otherwise stated, following (Malladi et al., 2023), half-precision training (F16) and mixed-precision training (FP16) are adopted by default on ZO and FO methods, respectively, to reduce the memory consumption. For FP16, the model is loaded with full precision while the training is carried out in half-precision, whereas F16 means both the model and training are in float16. More implementation details can be found in Appx. C.2, and their influence on memory efficiency is discussed in Sec. 4.3.
@@ 4.2 Experiment Results
Third, The performance of ZO methods exhibits high variance, as evidenced by the fluctuating relative rankings across different experimental scenarios, although extensive hyper-parameter search efforts have been made. For example, the effectiveness of ZO-Adam degrades dramatically in the (OPT-1.3B, Prompt) setting. In addition, the MeZO method (i.e., ZO-SGD) used in (Malladi et al., 2023) does not always emerge as the top-performing ZO optimizer for LLM fine-tuning across various settings. This is not surprising and could be largely attributed to the high variance of RGE (Nesterov & Spokoiny, 2017; Duchi et al., 2015).
@@
First, compared to the previous results, the performance gap among different ZO methods is much enlarged. In the meantime, the performance gap between FO and ZO methods is also widened. For example, in the experiments with WinoGrande, the FO methods (FO-SGD and FO-Adam) outperform all the other ZO methods by a large margin. This observation shows the scalability bottleneck intrinsic to ZO methods, when dealing with larger models and/or more complex tasks.
Second, certain ZO methods exhibit exceptional stability across varied conditions: Despite a general trend towards variability, specific ZO methods, e.g., ZO-Adam and ZO-SGD-MMT, demonstrate stability in their performance. This could be because these algorithms integrate variance-reduced optimization techniques (such as momentum and adaptive learning rate) to ZO optimization and become more adaptable and resilient to the variances of ZO gradient estimation (Chen et al., 2019).
Third, the LoRA tuning is consistently robust when paired with various ZO algorithms. This resilience across different ZO methods suggests that LoRA’s mechanism is inherently more adaptable to various ZO optimization strategies, providing a stable and reliable tuning approach in diverse settings. We will peer into the performance of LoRA below.
@@
In Tab. 3, we present how different ZO methods perform on (LoRA, COPA) and (LoRA, WinoGrande) across a wide range of LLM families. For ease of computation, we focus on a subset of ZO optimization methods, including ZO-SGD, ZO-SGD-CONS, and ZO-Adam. As we can see, in some scenarios with the COPA dataset, some BP-free methods exhibit effectiveness comparable to, or even superior to, that of FO methods (FO-SGD and FO-Adam). For example, Forward-Grad and ZO-Adam outperform the best FO method on model OPT-13B and Vicuna-7B. Conversely, for the more difficult task WinoGrande, a performance gap of $5\%\sim 6\%$ between FO and ZO methods still exists across different models.
Efficiency analysis. In Tab. 4, we present a comparison of the efficiency performance of various ZO/FO optimizers when fine-tuning the full OPT-13B model on the MultiRC dataset with a batch size of $4$. We evaluate the efficiency in the following dimensions: memory cost (in GB), the consumption of GPU resources (number of GPUs), and runtime cost per optimization iteration (in seconds). All the experiments are carried out in the same environment. First, from a memory efficiency standpoint, ZO-SGD, ZO-SGD-Cons, and ZO-SGD-Sign exhibit similar levels of efficiency and only necessitate a single GPU (A100) for LLM fine-tuning This is not surprising since these ZO methods employ relatively simple optimization steps and primarily rely on the utilization of RGE. Second, Forward-Grad seems to be the threshold beyond which the ZO optimization methods lose their memory efficiency advantage over FO methods, as seen with ZO-Adam, for example. Third, compared to FO methods, ZO optimization reduces runtime costs per iteration by $\sim 41.9\%$ for ZO-SGD versus FO-SGD.
Table 4: The peak memory cost (in GB), the required GPU resources, and the runtime cost (in seconds) of each optimizer when fine-tuning the full OPT-13B model on MultiRC with an averaged 400 context length. The order of included optimizers is ranked based on the memory cost. The per-iteration runtime in seconds (s) is averaged over 100 iterations.
@@
ZO-SGD $\mathbf{64}$ GB $\mathbf{1}\times$A100 $\mathbf{11.1}$s
ZO-SGD-Cons $\mathbf{64}$ GB $\mathbf{1}\times$A100 $27.6$s
ZO-SGD-Sign $\mathbf{64}$ GB $\mathbf{1}\times$A100 $\mathbf{11.1}$s
ZO-SGD-MMT $100$ GB 2×2$\times$2 ×A100 $\mathbf{11.1}$s
Forward-Grad $134$ GB 2×2$\times$2 ×A100 $16.7$s
FO-SGD $148$ GB 2×2$\times$2 ×A100 $19.1$s
ZO-Adam $158$ GB 2×2$\times$2 ×A100 $\mathbf{11.5}$s
FO-Adam $245$ GB 4×4$\times$4 ×A100 $19.4$s
Ablation study on query budget. Recall from (1) that Forward-Grad provides us with an unbiased gradient estimate with respect to the FO gradient, in contrast to the function value-based biased ZO gradient estimate. However, in the above experiments, we have not observed a significant advantage brought by Forward-Grad. We hypothesize that this is due to the smallest query budget $q=1$ we used, which, despite its query efficiency, may not fully showcase the unbiased benefit. Inspired by the above, Fig. 2 explores the impact of varying query budget ($q$) on the performance of Forward-Grad and ZO-SGD (i.e., MeZO). As we can see, both the accuracies of Forward-Grad and ZO-SGD are improved with an increased query budget. However, the performance rise for Forward-Grad is much more evident. For example, when the query number is larger than $500$, Forward-Grad outperforms ZO-SGD by a large margin $1\%\sim 2\%$ and approaches FO-SGD. These observations underscore that the inherent advantages of Forward-Grad become apparent only when a sufficient number of queries are used. The downside of using a higher query budget is a higher computation cost, which scales linearly with $q$.